OpenStudy (sparklestaraa):

A coffee shop pays Coffee Delivery Company A, a certain price for each disposable cup it orders plus a weekly delivery charge to remain on the driver’s delivery route. The cups are purchased in increments of 500. To quickly determine how much the coffee shop will be spending on cups before their arrival, the owner created the following table:

OpenStudy (sparklestaraa):

A) The price of the cups, p(x), is a function of the number of x cups ordered. Using the table, determine the average rate of change for the first 1,000 cups ordered and then for all 3,000 cups ordered. What does this tell you about the function? B) Use the table to evaluate and interpret p(0). What is a possible explanation for this? C ) Sketch a graph, labeling its key features, to show the price the coffee shop would pay Coffee Delivery A to have between 0 and 3,000 cups delivered each week. D) Create a model using function notation that represents how the two quantities, cups and cost, are related. E. The coffee shop found another delivery company that sells orders at increments of 500 cups, Coffee Delivery B. They charge $3.50 each week to be on their delivery route and charge 3.9 cents per disposable cup. -Make a function using the information about the second delivery company. -Graph the price the coffee shop would pay for Coffee Delivery B to deliver the same amount of cups on the same graph from Part C, using a different color for the new line. -If the coffee shop can change delivery companies every three months, when should they consider Coffee Delivery A, and when should they consider Coffee Delivery B?

OpenStudy (sparklestaraa):

@CoconutJJ will give medal I'm desperate for help

OpenStudy (sparklestaraa):

@wio

OpenStudy (sparklestaraa):

@ganeshie8

OpenStudy (coconutjj):

What are you confused on?

OpenStudy (sparklestaraa):

everything tbh

OpenStudy (coconutjj):

Lets try A first

OpenStudy (coconutjj):

How do you calculate Avg Rate of Change

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (sparklestaraa):

@fatZak

OpenStudy (sparklestaraa):

@DanJS last thing :D then my linear lessons are over :D

OpenStudy (danjs):

A) The price of the cups, p(x), is a function of the number of x cups ordered. Using the table, determine the average rate of change for the first 1,000 cups ordered and then for all 3,000 cups ordered. What does this tell you about the function?

OpenStudy (danjs):

Average rate of change is = slope of secant line connecting the 2 points \[ave~change~=\frac{ p(1000) - p(0) }{ 1000 - 0 } = \frac{ $44 - $7 }{ 1000 } = \frac{ $37 }{ 1000 }\]

OpenStudy (danjs):

change to decimal if you want

OpenStudy (danjs):

Average for 3000 cups ordered --- (3000,118) and (0 , 7) are the 2 points to use for the secant line. The slope of the secant is... \[ave~change~=\frac{ 118 - 7 }{ 3000 } = \frac{ 111 }{ 3000 }\]

OpenStudy (danjs):

p(x) is the price for x cups ordered, p(0), means he orders zero cups, and p(0)=$7, he is charged 7 dollars for ordering zero cups. That must be the charge for shipping, extra costs

OpenStudy (sparklestaraa):

this is all for A?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

That was A and B

OpenStudy (danjs):

All of these long story probs are exactly the same, just a different story. A graph of a line.

OpenStudy (sparklestaraa):

Yes ur right if u write c and d can u write the letter before the explanation

OpenStudy (sparklestaraa):

which is B? @DanJS

OpenStudy (sparklestaraa):

is this b p(x) is the price for x cups ordered, p(0), means he orders zero cups, and p(0)=$7, he is charged 7 dollars for ordering zero cups. That must be the charge for shipping, extra costs

OpenStudy (danjs):

B) interpret p(0). p(0), means he orders zero cups, and p(0)=$7, he is charged 7 dollars for ordering zero cups. That must be the charge for shipping, extra costs

OpenStudy (danjs):

C) to sketch the graph, label the x axis cups ordered, and the Y axis, dollar paid plot each of those points ( cups, dollars paid)

OpenStudy (danjs):

The slope of the line in the graph will be 18.5 dollars per 500 cups

OpenStudy (sparklestaraa):

can u make the graph with desmos like i said last night i dont have graph paper or anything and dont know how to put it in to get the graph

OpenStudy (danjs):

D) 18.5 dollars / 500 cups = 0.037 dollars / 1cup, that is the slope p(x) = slope*(x) + y intercept p(x) = 0.037*x + 7

OpenStudy (danjs):

Just go to desmos, and type in y = 0.037*x+7 make the x axis from 0 to 3000 y axis from 0 to like 150 dollars

OpenStudy (danjs):

E. The coffee shop found another delivery company that sells orders at increments of 500 cups, Coffee Delivery B. They charge $3.50 each week to be on their delivery route and charge 3.9 cents per disposable cup. -Make a function using the information about the second delivery company. c(x) = cost of purchasing x cups @ $0.039 per cup, and 3.50 fee c(x) = 0.039*x + 3.50

OpenStudy (danjs):

add y = 0.039x + 3.5 to the graph of the first part

OpenStudy (sparklestaraa):

add that to the graphing calculator?

OpenStudy (sparklestaraa):

how do i make the x and y axis with #'s in desmos?

OpenStudy (sparklestaraa):

its still not working for me @DanJS

OpenStudy (danjs):

Blue Line is First Part, Has Y Intercept of 7 Dollars for Zero Cups Red Line is Second Part, Has Y intercept of 3.50 dollars for zero cups.

OpenStudy (danjs):

Hit the wrench button in the top right to change the axis ranges

OpenStudy (danjs):

X axis - # Cups Y axis - Cost

OpenStudy (sparklestaraa):

:D thankssss

OpenStudy (danjs):

not sure if it is a question, but the point in the middle where they cross, is where they both will be the same price. For 1750 cups , they both will cost 71.75

OpenStudy (danjs):

For the very last thing, not sure -If the coffee shop can change delivery companies every three months, when should they consider Coffee Delivery A, and when should they consider Coffee Delivery B?

OpenStudy (danjs):

3 months , would be i guess, 12 weeks

OpenStudy (danjs):

For the first company, 12 weeks would cost a fee of 12*$7 = 84 dollars second company 12 weeks would cost fee 12 * 3.5 = 42 dollars

OpenStudy (danjs):

but the first company costs less per cup bought, 37 cents vs the second company 39 cents per cup

OpenStudy (danjs):

that's all i got

OpenStudy (danjs):

lol you have by far the longest probs on here

OpenStudy (sparklestaraa):

xD yayy thanks so much :D

OpenStudy (sparklestaraa):

oh oh oh and the graph for C is that the same one in Blue Line is First Part, Has Y Intercept of 7 Dollars for Zero Cups Red Line is Second Part, Has Y intercept of 3.50 dollars for zero cups.

OpenStudy (danjs):

yeah

OpenStudy (sparklestaraa):

:D thankssss :D ur the best really and I'm going to write a testimonial for u :D

OpenStudy (danjs):

haha, thanks :), no more 20 page probs please!

OpenStudy (danjs):

nah its cool

OpenStudy (sparklestaraa):

xD hhhhhh ok i promise well i cant keep that promise XD but i'll give u a page at a time xD

OpenStudy (danjs):

Just the all are the same, just different story to go with them. Equations for a line.